OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

@AlexandervonHumboldt2

Nnesha (nnesha):

use same formula to find slope first :)

OpenStudy (alexandervonhumboldt2):

(2+4)/(6-8)=6/-2=-3 y=-3x+b subs points 2=-3*6+b

OpenStudy (alexandervonhumboldt2):

b=20 y=-3x+20

OpenStudy (alexandervonhumboldt2):

3x-y=20

OpenStudy (anonymous):

thnx

Nnesha (nnesha):

i hope i will not see this again last warning LAST WARNING